A student has 6.07 grams of rubidium and an unlimited supply of both carbon and nitrogen. Determine the maximum mass of the fictional compound Rb4C7N5 that can be created.


1
Expert's answer
2021-04-12T00:04:40-0400

Molar Mass of Rb = 85.4678g/mol

= 85.4678×4 = 341.8712g

Molar Mass of carbon = 12.0107

12.0107×7 = 84.0749g

Molar Mass of Nitrogen = 14.0067

= 14.0067 × 5 = 70.0335

= (6.07/85.4678)×341.8712 = 24.28g

= 24.28+84.0749+70.0335

= 178.3884g